# Exploring a New Design Paradigm for Omnidirectional MAVs for Minimal Actuation and Internal Force Elimination: Theoretical Framework and Control

### v2

@@ Abstract
This paper presents a novel concept for achieving omnidirectionality in a multirotor aerial vehicle (MAV) that uses only 6 inputs and ensures no internal forces at the equilibria. The concept integrates a single actively-tilting propeller along with 3 pendulum-like links, each carrying a propeller, connected by passive universal joints to the main body. We show that this design ensures omnidirectionality while minimizing the internal forces and without resorting to overactuation (i.e., more than 6 inputs). A detailed dynamic model of the multi-link MAV is first developed. Afterwards, the analysis identifies the equilibrium configurations and illustrates that a forced equilibrium exists for every pose of the MAV’s main platform. In order to render this equilibrium asymptotically stable for the closed-loop system, a coordinate-invariant nonlinear controller is constructed using dynamic feedback linearization and backstepping techniques with the main platform configuration error being the left-trivialized error on $SE(3)$. The stability of the closed-loop system is then investigated by employing standard Lyapunov arguments on the zero dynamics. We conclude by providing numerical Gazebo simulations, whose videos are available at https://youtu.be/SWIt0CSwNUo, validating our approach. They demonstrate the MAV capability to perform decoupled attitude and translational motions under parametric uncertainty and actuators noise.
Nomenclature
$SE(3)$
Special Euclidean group of rigid-body transformations.
se(3){se}(3)
Lie algebra of $SE(3)$.
$\mathbb{T}^{n}$
$n$-torus manifold (a compact Lie group).
$C^{0}_{j}\in SE(3)$
Configuration of $B_{j}$’s frame $\mathcal{F}_{p_{j}}$ in the world frame $\mathcal{F}_{w}$
$V^{0}_{1}\in\mathbb{R}^{6}$
Right-invariant base twist expressed in, and relative to, $\mathcal{F}_{w}$
[]:[*]: $\mathbb{R}^{6}\to se(3)$ An isomorphism with an inverse $\hat{*}$.
$(*)_{\times}$ A skew-symmetric matrix of vector $\in\mathbb{R}^{3}$.
$J_{1j}$
Universal joint connecting the base to link $B_{j}$.
$S_{j}\in\mathbb{R}^{6\times 2}$
$J_{1j}$’s spatial instantaneous screw axis, $S_{j}=(\,S_{j1}\;S_{j2}\,)$.
$S_{p}=\operatorname{blkdiag}(S_{j})$
Block-diagonal matrix of screws $S_{j}$.
$\operatorname{Ad}_{C}:$
$SE(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $SE(3)$.
$\operatorname{ad}_{V}:$
$se(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $se(3)$.
$R_{k}(*)$
Pure rotation about axis $k$ by a angle *.
$e_{i}$
zero vector with one in index $i$.
$M$
Generalized mass (inertia) matrix.
$M_{bb},\ M_{bp},\ M_{pp}$
Base, coupling, and joint sub-matrices of $M$.
$h$
Centrifugal and Coriolis terms.
$g_{b},\,g_{p}$
Gravitational terms of base and joints, $g=(\,g_{b}\,\,g_{p}\,)^{T}$.
$W^{0}_{\mathrm{app,b}}$
Applied wrench on base’s CoM expressed in $\mathcal{F}_{w}$.
$W^{0}_{\mathrm{app,p}}$
Applied wrench on links $B_{2}$, $B_{3}$, $B_{4}$’s CoM expressed in $\mathcal{F}_{w}$ and projected onto the respective joints through the corresponding screw in $S_{p}$.
$\tau_{dr_{i}}$
Drag torque about propeller $i$’s axis.
$\tau_{f_{ji}}$
Viscous friction at joint $q_{ji}$’s axis. $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{ji}})\big)^{T}$
$\tau_{s_{j}}$
Torsion spring torque applied at joint $q_{j1}$’s axis.
$\tau_{dj}$
Propeller $j$’s drag torque factorized on $q_{j1}$’s axis at equilibrium.
$m_{b},m_{j}$
Mass of base and link $B_{j}$.
$c$
Length from joints axes intersection point to $B_{j}$’s CoM.
$k_{i}$
Drag-to-thrust ratio of propeller $i$.
$\mathrm{PJD}_{1},\ \mathrm{PJD}_{2}$
Passive-joint designs type 1 and type 2.
$P_{se(3)}(A)$
projects a matrix $A\in\mathbb{R}^{4\times 4}$ on $se(3)$.
$_{e}$
Variable * evaluated at equilibrium.
$^{[i]}$
Value of variable * pertaining to joint design PJDi.
@@
I-A Research Problems and Related Work
@@
When designing a new MAV, one can easily be convinced that it is desirable to obtain all the angular velocities of the propellers aligned with each other and with the gravity force at any static hovering equilibrium: this avoids internal propeller-generated forces canceling each other at the cost of a higher energy consumption; moreover, it is desirable to keep the number of actuators to the minimum (equal to the task dimension), to reduce the payload (which, in turn, affects the vehicle’s endurance), the cost, the maintenance requirements, etc. Furthemore, for tasks requiring high dexterity like the ones mentioned before, it is desirable for the MAV to be able to statically hover at any pose (in SE(3) in general). Last but not least, the design of a MAV must also allow the controlled stabilization of its hovering equilibria. We label these four desirable properties as the Internal-force Property, the Input-dimension Property, the Equilibria Property, and the Stabilizability Property, respectively.
@@
When restricting the problem to the vertical 2D space, we demonstrated in [6] that a MAV design with all the above four properties exists. Especially, it can be stabilized at any static equilibrium pose in SE(2) with propellers parallel to each other and the gravity vector, and it has only three actuators. The actuators are two brushless motors attached to one propeller each and one servomotor that actively tilts one of the two propellers, while the other one is free to passively rotate and has a center of mass not coincident with the rotation point. However, extending such a result to the 3D space is very far from being trivial and is an open problem. Just to mention two obstacles to such extension: SE(3) is a fundamentally different manifold than SE(2) and the dynamic effects related to single revolute joints in 2D cannot be extended to the case of multiple joints in 3D, see model in Sec. II-B. Moreover, in the 3D space, the propeller drag torques arise and have to be taken into account, see discussion in Sec. III-A.
@@
We will now revise the main MAV designs proposed in the literature and show that none of them has all the four desirable properties described above.
@@
Standard underactuated MAVs like quadrotors and hexarotors tick the Internal-force Property as they have all propellers fixedly oriented with spinning axes parallel to each other; however, because of the fixed attachment between the propellers and the body frame, their only hovering equilibria are characterized by the attitudes where the $z$ axis is aligned with the gravity vector. Thus, they do not tick the Equilibria property.
@@
Aerial vehicles equipped with the ability to hover while maintaining a larger set of orientations, known as fully-actuated MAVs, are surveyed in [7]. These MAVs cannot be stabilized at any pose but at a subset of poses in SE(3) determined by the thrust limits, the design kinematics, and joint limits. Such subset contains any position but only a subset of all the orientations in general. A subclass of this type of MAVs, called omnidirectional MAV [8], can hover at any pose: namely, they have the Equilibria property mentioned above. These aerial vehicles can be divided in two classes.
@@
The first class of designs that have been proposed to achieve full actuation or omnidirectionality are realized by fixedly tilting the propellers in such a way that their spinning axes are not all parallel to each other. As a consequence, all of these designs do not have the Internal-force Property. Examples are [9], a fully actuated hexarotor, and [10, 11], omnidirectional hexarotors. Other omnidirectional fixedly-tilted-propeller designs do not comply with the Input-dimension Property, either; they are, for example, the hepta-rotor presented in [12] or the octarotors proposed in [13, 14].
@@
I-B Contributions
The main contribution of this work is to propose the first MAV design, to the best of the authors’ knowledge, having the following four properties in SE(3):
@@ I-B Contributions
These properties capture major requirements for enabling energy-efficient omnidirectional flight. In particular, the ability to hover at any attitude in $SO(3)$ is the defining feature of omnidirectional MAVs [14, 15], hence expanding their reachable workspace and manipulation capabilities [12]. Ensuring stabilizability of these equilibria is essential for any practical deployment. Moreover, achieving full 6-DOF control with only 6 actuators alongside eliminating internal forces at hovering directly contributes to significantly improved energy efficiency, simplified input allocation, and extended flight endurance [21, 13, 22]. Together, these properties illustrate the impact of the proposed MAV concept within the field of aerial robotics.
I-C Motivations Behind the Design Idea
@@
I-D Outline
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For ease of exposition, all the formal proofs, derivations, and explicit calculation formulas are collected in appendices of [29].
II Proposed Design and Dynamical Model
Figure 1: Conceptual representation of the proposed MAV idea. The origins of the frames $\mathcal{F}_{b}$ and $\mathcal{F}_{p_{2}}$ are attached to CoMs of the base $B_{1}$ and the propeller link $B_{2}$, respectively. The shown configuration is when $\mathcal{F}_{b}$ coincides with $\mathcal{F}_{w}$ and all the frames of the links have the same orientation as $\mathcal{F}_{w}$. The joint $q_{j1}$ and $q_{j2}$ positions are zero when $\mathcal{F}_{p_{j}}$ has the same orientation as $\mathcal{F}_{b}$, $j=\{2,\ldots,5\}$. The link $B_{5}$ is attached to $B_{1}$ through an active joint, unlike the the rest of the joints which are passive.
II-A Mechanical Design Idea
@@
The vehicle concept, shown in Fig. 1, consists of a main rigid body, referred to as the base and denoted by $B_{1}$, and four smaller bodies, referred to as links and denoted by $B_{2},B_{3},B_{4}$, and $B_{5}$. Each link carries a rigidly-attached motor driving a propeller and is connected to the base by two revolute joints. The propeller, the motor, the link, and the two revolute joints together form an actuation unit. For each unit, the joints’ mutually-orthogonal rotation axes intersect at a point, called the center of rotation of the link, coinciding with the propeller center. The actuation units are mounted on $B_{1}$ such that all centers of rotation lie on the same plane, resulting in a coplanar design. The centers of mass (CoMs) of $B_{2},B_{3}$, and $B_{4}$ are located below their respective centers of rotation, achieved by placing the motors sufficiently low. This makes these links behave as spatial pendula, with the joints acting as pivots. In contrast, $B_{5}$’s CoM coincides with its center of rotation. $B_{1}$ is connected to $B_{5}$ by an actuated universal joint with configuration $q_{a}=\operatorname{col}(q_{5i})\in\mathbb{T}^{2}$, and to $B_{2},B_{3}$, and $B_{4}$ by passive universal joints with configurations $q_{p}=\operatorname{col}(q_{ji})\in\mathbb{T}^{6}$, $j=2,\ldots,5$, $i=1,2$.
@@ III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
The state equilibria set for PJDi is denoted by $\mathbb{D}^{[i]}_{x_{e}}$ whereas $\mathbb{D}^{[i]}_{f_{e}}$ signifies their associated thrust equilibrium sets. The explicit relations are given in Sec. -D4 [29].
@@ IV Stabilization of the Equilibria
Figure 3: Block diagram of the proposed control scheme. It takes a desired base pose $C^{0}_{1e}\in SE(3)$ and converts it into a reference base twist $\zeta_{1e}(t)$ through configuration error feedback $E_{c}$. Dynamic FBL is applied to the twist dynamics, which are stabilized using a PID controller. The accelerations are efficiently computed via a modified Articulated-Body Inertia Algorithm (ABA) in Sec. -B of [29].
Theorem 1.
@@
Proposition 2.
@@
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in\mathbb{R}^{6}$ are respectively given in (86) and (98) [29] and shown in Fig. 3. $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
@@
We note that the system (1) evolves on the product Lie group $SE(3)\times\mathbb{T}^{8}$ due to the internal joint dynamics. This motivates our use of the normal-form approach [27], as it facilitates the study of the associated zero dynamics. Although the proposed controller, shown in Fig. 3, employs tools often considered coordinate-dependent, the Lie-group formulation guarantees that the control laws are coordinate-invariant, i.e. their validity does not depend on a specific local parameterization of the configuration $SE(3)\times\mathbb{T}^{8}$.

### v3

# The PenduMAV: A Six-Input Omnidirectional MAV without Internal Forces - Design, Dynamics, and SE(3) Control

@@ Abstract
We introduce the PenduMAV, an exactly actuated (6-input) omnidirectional multirotor that structurally eliminates internal forces at equilibria. The vehicle features one actively tilting propeller and three propellers mounted on passive pendulum links via universal joints. This architecture achieves full 6D wrench generation while avoiding the structural and energetic costs of input redundancy and internal forces. After deriving the full multibody dynamics, we demonstrate that a forced equilibrium exists for every main platform pose. To asymptotically stabilize the closed-loop system, we design a coordinate-invariant nonlinear controller based on dynamic feedback linearization and backstepping, utilizing the left-trivialized error on $SE(3)$. System stability is formally guaranteed through Lyapunov analysis of the zero dynamics. Finally, Gazebo simulations (videos available at https://www.youtube.com/playlist?list=PL4N8pJgvqASQX6AWEpg3NCZ6QdGBPfbXq) validate the approach, showcasing fully decoupled attitude and translational tracking under parametric uncertainty and actuator noise.
@@
Research Problems and Related Work
@@
When designing a new MAV, one can easily be convinced that it is desirable to obtain all the angular velocities of the propellers aligned with each other and with the gravity force at any static hovering equilibrium: this avoids internal propeller-generated forces canceling each other at the cost of a higher energy consumption; moreover, it is desirable to keep the number of actuators to the minimum (equal to the task dimension), to reduce the weight (which, in turn, affects the vehicle’s endurance), the cost, the maintenance requirements, etc. Furthermore, for tasks requiring high dexterity like the ones mentioned before, it is desirable for the MAV to be able to statically hover at any pose (in SE(3) in general). Last but not least, the design of a MAV must also allow the controlled stabilization of its hovering equilibria. We label these four desirable properties as the Internal-force Property, the Input-dimension Property, the Equilibria Property, and the Stabilizability Property, respectively.
@@
When restricting the problem to the vertical 2D space, we demonstrated in [6] that a MAV design with all the above four properties exists. Especially, it can be stabilized at any static equilibrium pose in SE(2) with propellers parallel to each other and the gravity vector, and it has only three actuators. The actuators are two brushless motors attached to one propeller each and one servomotor that actively tilts one of the two propellers, while the other one is free to passively rotate and has a center of mass not coincident with the rotation point. However, extending such a result to the 3D space is very far from being trivial and is an open problem. Just to mention two obstacles to such extension: SE(3) is a fundamentally different manifold than SE(2) and the dynamic effects related to single revolute joints in 2D cannot be extended to the case of multiple joints in 3D, see model in Sec. II-C. Moreover, in the 3D space, the propeller drag torques arise and have to be taken into account, see discussion in Sec. III-A.
State of the Art
@@
We revise the main MAV designs proposed in the literature and show that none of them has all the four desirable properties described above.
@@
Standard underactuated MAVs like quadrotors and hexarotors tick the Internal-force Property as they have all propellers fixedly oriented with spinning axes parallel to each other; however, because of the fixed attachment between the propellers and the body frame, the only hovering equilibria are characterized by the attitudes where the $z$ axis is aligned with gravity. Thus, they do not tick the Equilibria property.
@@
MAVs equipped with the ability to hover while maintaining a larger set of orientations, known as fully-actuated MAVs, are surveyed in [7]. These MAVs cannot be stabilized at any pose but at a subset of poses in SE(3) determined by the thrust limits, the design kinematics, and joint limits. Such subset contains any position but only a subset of all the orientations in general. A subclass of this type of MAVs, called omnidirectional MAV [8], can hover at any pose: namely, they have the Equilibria property mentioned above. These aerial vehicles can be divided in two classes.
@@
The first class of designs that have been proposed to achieve full actuation or omnidirectionality are realized by fixedly tilting the propellers in such a way that their spinning axes are not all parallel to each other. As a consequence, all of these designs do not have the Internal-force Property. Examples are [9], a fully actuated hexarotor, and [10, 11], omnidirectional hexarotors. Other omnidirectional fixedly-tilted-propeller designs do not comply with the Input-dimension Property, either; they are, for example, the hepta-rotor of [12] or the octarotors proposed in [13, 14].
@@
Contributions
The main contribution of this work is to propose the PenduMAV concept, the first MAV design, to the best of the authors’ knowledge, having the following four properties:
@@
These properties capture major requirements for enabling energy-efficient omnidirectional flight. In particular, the ability to hover at any attitude in $SO(3)$ is the defining feature of omnidirectional MAVs [14, 15], hence expanding their reachable workspace and manipulation capabilities [12]. Ensuring stabilizability of these equilibria is essential for any practical deployment. Moreover, achieving full 6-DOF control with only 6 actuators alongside eliminating internal forces at hovering directly contributes to significantly improved energy efficiency and extended flight endurance [21, 13, 22]. Together, these properties illustrate the impact of the proposed MAV concept within the field of aerial robotics.
The rest of the paper is organized as follows. Sec. II describes the dynamic model of the proposed novel MAV conceptual idea, whereas Sec. III identifies the equilibria sets of that model being all the poses in SE(3) and all having the internal force property. Afterwards, we investigate and demonstrate that the system can be made asymptotically stable at all these equilibria by a state feedback in Sec. IV. Numerical simulations where the MAV performs decoupled motions are the subject of Sec. V. We conclude by giving some remarks on the future work in Sec. VI. For a streamlined presentation, the formal proofs, derivations, and explicit calculation formulas are collected in the appendices of [29].
TABLE I: Nomenclature
Symbol Description
$SE(3)$ Special Euclidean group of rigid-body transformations.
se(3){se}(3) Lie algebra of $SE(3)$.
$\mathbb{T}^{n}$ $n$-torus manifold (a compact Lie group).
$C^{0}_{j}\in SE(3)$ Configuration of $B_{j}$’s frame $\mathcal{F}_{p_{j}}$ in the world frame $\mathcal{F}_{w}$.
$V^{0}_{1}\in\mathbb{R}^{6}$ Right-invariant base twist expressed in, and relative to, $\mathcal{F}_{w}$.
[][*] $\mathbb{R}^{6}\to se(3)$ An isomorphism with an inverse $\hat{*}$.
$(*)_{\times}$ The skew-symmetric matrix of the vector $\in\mathbb{R}^{3}$.
$J_{1j}$ Universal joint connecting the base to link $B_{j}$.
$S_{j}\in\mathbb{R}^{6\times 2}$ $J_{1j}$’s spatial instantaneous screw axis, $S_{j}=(\,S_{j1}\;S_{j2}\,)$.
$S_{p}$ Block-diagonal matrix of screws $S_{j}$, $\operatorname{blkdiag}(S_{j})$.
$\operatorname{Ad}_{C}$ $SE(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $SE(3)$.
$\operatorname{ad}_{V}$ $se(3)\times se(3)\rightarrow se(3)$ Adjoint representation of $se(3)$.
$R_{k}(*)$ Pure rotation about axis $k$ by a angle *.
$e_{i}$ Zero vector with a one at index $i$.
$M$ Generalized mass (inertia) matrix.
$M_{bb},M_{bp},M_{pp}$ Base, coupling, and joint sub-matrices of $M$.
$h$ Centrifugal and Coriolis terms.
$g_{b},g_{p}$ Gravitational terms of base and joints, $g=(\,g_{b}\,\,g_{p}\,)^{T}$.
$W^{0}_{\mathrm{app,b}}$ Applied wrench on base’s CoM expressed in $\mathcal{F}_{w}$.
$W^{0}_{\mathrm{app,p}}$ Applied wrench on links $B_{2},B_{3},B_{4}$’s CoM in $\mathcal{F}_{w}$ projected onto respective joints via $S_{p}$.
$\tau_{dr_{i}}$ Drag torque about propeller $i$’s axis.
$\tau_{f_{ji}}$ Viscous friction at joint $q_{ji}$’s axis. $\bar{\tau}_{f}=\big(0_{6\times 1}\;\;\operatorname{col}(\tau_{f_{ji}})\big)^{T}$.
$\tau_{s_{j}}$ Torsion spring torque applied at joint $q_{j1}$’s axis.
$\tau_{dj}$ Propeller $j$’s drag torque factorized on $q_{j1}$’s axis at equilibrium.
$m_{b},m_{j}$ Mass of base and link $B_{j}$.
$c$ Length from joints axes intersection point to $B_{j}$’s CoM.
$k_{i}$ Drag-to-thrust ratio of propeller $i$.
$\mathrm{PJD}_{1},\mathrm{PJD}_{2}$ Passive-joint designs of type 1 and type 2.
$P_{se(3)}(A)$ Projects a matrix $A\in\mathbb{R}^{4\times 4}$ onto $se(3)$.
$_{e}$ Variable * evaluated at equilibrium.
$^{[i]}$ Value of variable * pertaining to joint design PJDi.
II Proposed Design and Dynamical Model
II-A Motivations Behind the Design Idea
@@
II-B Mechanical Design Idea
Figure 1: Conceptual representation of the PenduMAV idea. The origins of the frames $\mathcal{F}_{b}$ and $\mathcal{F}_{p_{2}}$ are attached to CoMs of the base $B_{1}$ and the propeller link $B_{2}$, respectively. The shown configuration is when $\mathcal{F}_{b}$ coincides with $\mathcal{F}_{w}$ and all the frames of the links have the same orientation as $\mathcal{F}_{w}$. The joint $q_{j1}$ and $q_{j2}$ positions are zero when $\mathcal{F}_{p_{j}}$ has the same orientation as $\mathcal{F}_{b}$, $j=\{2,\ldots,5\}$. The link $B_{5}$ is attached to $B_{1}$ through an active joint, unlike the the rest of the joints which are passive.
@@ II-B Mechanical Design Idea
The PenduMAV concept, shown in Fig. 1, consists of a main rigid body, referred to as the base and denoted by $B_{1}$, and four smaller bodies, referred to as links and denoted by $B_{2},B_{3},B_{4}$, and $B_{5}$. Each link carries a rigidly-attached motor driving a propeller and is connected to the base by two revolute joints. The propeller, the motor, the link, and the two revolute joints together form an actuation unit. For each unit, the joints’ mutually-orthogonal rotation axes intersect at a point, called the center of rotation of the link, coinciding with the propeller center. The actuation units are mounted on $B_{1}$ such that all centers of rotation lie on the same plane, resulting in a coplanar design. The centers of mass (CoMs) of $B_{2},B_{3}$, and $B_{4}$ are located below their respective centers of rotation, achieved by placing the motors sufficiently low. This makes these links behave as spatial pendula, with the joints acting as pivots. In contrast, $B_{5}$’s CoM coincides with its center of rotation. $B_{1}$ is connected to $B_{5}$ by an actuated universal joint with configuration $q_{a}=\operatorname{col}(q_{5i})\in\mathbb{T}^{2}$, and to $B_{2},B_{3}$, and $B_{4}$ by passive universal joints with configurations $q_{p}=\operatorname{col}(q_{ji})\in\mathbb{T}^{6}$, $j=2,\ldots,5$, $i=1,2$.
@@ III-B Minimal Thrust at Equilibrium vs Presence of Internal Forces
The state equilibria set for PJDi is denoted by $\mathbb{D}^{[i]}_{x_{e}}$ whereas $\mathbb{D}^{[i]}_{f_{e}}$ signifies their associated thrust equilibrium sets. The explicit relations are given in Sec. -D4 of [29].
@@ IV Stabilization of the Equilibria
Figure 3: Block diagram of the control scheme. It takes a desired base pose $C^{0}_{1e}\in SE(3)$ and converts it into a reference base twist $\zeta_{1e}(t)$ through configuration error feedback $E_{c}$. Dynamic FBL is applied to the twist dynamics, which are stabilized using a PID controller. The accelerations are efficiently computed via a modified Articulated-Body Inertia Algorithm (ABA) in Sec. -B of [29].
Theorem 1 (Feedback Linearization Compensator and Zero Dynamics (inner loop)).
@@
Proposition 2 (Asymptotic Stabilization of Equilibria (outer loop)).
@@
where $\zeta_{1e}$, $\dot{\zeta}_{1e}$, $\ddot{\zeta}_{1e}$ $\in\mathbb{R}^{6}$ are respectively given in (86) and (98[29] and shown in Fig. 3. $\bar{K}_{d}$ and $\bar{K}_{a}$ $\in\mathbb{R}^{6\times 6}$ are positive definite diagonal gain matrices chosen such that to make (ABK)(A-BK) Hurwitz where
@@
We note that the system (1) evolves on the product Lie group $SE(3)\times\mathbb{T}^{8}$ due to the internal joint dynamics. This motivates our use of the normal-form approach [27], as it facilitates the study of the associated zero dynamics. Although the proposed controller, shown in Fig. 3, employs tools often considered coordinate-dependent, the Lie-group formulation guarantees that the control laws are coordinate-invariant, i.e., their validity does not depend on a specific local parameterization of the configuration $SE(3)\times\mathbb{T}^{8}$.